A stone falls from a railroad overpass which is 36 feet high into the path of a train which is approaching the overpass with uniform speed. If the stone falls when the train is 30 feet away from the overpass and the stone hits the ground just as the train arrives at that spot. How fast is the train moving?
1
Expert's answer
2020-09-15T10:04:18-0400
"t=\frac{s}{v}=\sqrt{\frac{2h}{g}}\\\frac{30}{v}=\sqrt{\frac{2(36)}{32.174}}\\v=20\frac{ft}{s}"